OpenStudy (anonymous):

Find the lateral area of a pyramid which has 3 and 9 as dimensions to the nearest whole unit.

OpenStudy (anonymous):

its 352m^2

Directrix (directrix):

LA of regular pyramid = 1/2 * p * L where p is the perimeter of the base and L is the slant height LA = 1/2* 8*4* 22 LA = ?

OpenStudy (anonymous):

its 352m^2

OpenStudy (anonymous):

Is it really?

OpenStudy (anonymous):

It might be...I'm not sure.

OpenStudy (anonymous):

it is... im a professor

Directrix (directrix):

I am checking this formula.

OpenStudy (anonymous):

Okay. I trust both of you. If you are both saying it is 352, then it is probably right.

Directrix (directrix):

Formula is okay. I concur with LL.

OpenStudy (anonymous):

I am going to close this question and post a new one. I only have 2 more questions thats it.

OpenStudy (anonymous):

Thank you

Directrix (directrix):

We need to revisit that cylinder problem from last night. Maybe LL will see the error.

OpenStudy (anonymous):

Who is LL?